Jim ,Carla and Tommy are members of the same family.Carla is 5 years older than Jim.Tommy is 6 years older than Clara. The sum of their 3 ages is 31 years. How old is each one of them?
Answer by macston(5194)

.
J=Jim's age; C=Carla's age=J+5yrs; T=Tommy's age=C+6yrs
.
J+C+T=31yrs
J+C+(C+6yrs)=31yrs
J+(J+5yrs)+((J+5yrs)+6yrs)=31yrs
3J+16yrs=31yrs
3J=15yrs
J=5yrs
C=J+5yrs=5yrs+5yrs=10yrs
T=C+6yrs=10yrs+6yrs=16yrs
.
ANSWER: Jim is 5 years, Carla is 10 yrs, and Tommy is 16 yrs.
.
CHECK:
.
J+C+T=31yrs
5yrs+10yrs+16yrs=31yrs
31yrs=31yrs